A fruit company delivers its fruit in two types of boxes: large and small. A delivery of 4 large boxes and 8 small boxes has a total weight of 197 kilograms. A delivery of 2 large boxes and 3 small boxes has a total weight of 83 kilograms. How much does each type of box weigh?
Answer by ikleyn(53763)

.

Let x = the weight of the large box;  y be the weight of the small box.

Then from the condition, you have these two equations


    4x + 8y = 197     (1)

    2x + 3y = 83      (2)


Multiply equation (2) by 2 (both sides).  Keep equation (1) as is

    4x + 8y = 197     (1')

    4x + 6y = 166     (2')


Subtract equation (2') from (1').  You will get

         2y = 197 - 166 = 31;

          y = 31/2 = 15.5.


Then from equation (2)

         2x = 83 - 3*15.5 = 36.5

          x               = 36.5/2 = 18.25,


ANSWER.  The large box weights 18.25 kg;  the small box weights 15.5 kg.

Solved.

---------------

You use very old source, where weight is measured in kilograms.

This practice was stopped in 60-es last century, when after changing the measurement system,
kilograms were replaced by newtons as the unit of force/weight.